At the Malta ferry, ships leave every 20 minutes and boats leave every 15 minutes. the first ship and the first boat leave at 6.am . at what time wiil the ship and boat next leave together
Found 2 solutions by fractalier, Alan3354:
Answer by fractalier(6550)

The next time is found by looking at the least common multiple of 15 and 20.
15,30,45,60,75,90, etc. and
20,40,60,80,100, etc.
You can see that at 60 minutes they will be leaving at the same time.
Thus, the answer is 7 am.

Answer by Alan3354(69443)

At the Malta ferry, ships leave every 20 minutes and boats leave every 15 minutes. the first ship and the first boat leave at 6.am . at what time wiil the ship and boat next leave together
================
Find the LCM (Least Common Multiple) of 15 and 20
4*15 = 3*20 = 60
LCM = 60 minutes
--> 7 AM